A person walks 25.00 north of east for 3.10 km. How far would she have to walk due north and due east to arrive at the same location?
1
Expert's answer
2020-10-13T10:41:21-0400

Let's draw a drawing for the task

Where

"OB=3.10km"

Then

a person will cover the distance to the north

"OA=OB \cdot \sin(25^0)=3.1 \cdot 0.423=1.31 km"

a man will walk the distance to the east

"AB=OB \cdot \cos(25^0)=3.1 \cdot 0.906=2.81 km"